OpenStudy (anonymous):

simplify this sqrt 32 x^7

OpenStudy (luigi0210):

see we can't really do anything with that 7..

OpenStudy (anonymous):

sure you can

OpenStudy (anonymous):

i have this sqrt 32 x^7=4rad2x^7=4 sqrt 2x^7 i just have to simplify the 7

OpenStudy (anonymous):

two goes in to 7 3 times with a remainder of 1 so \[\sqrt{x^7}=x^3\sqrt{x}\]

OpenStudy (luigi0210):

My mind if blown

OpenStudy (luigi0210):

*is

OpenStudy (anonymous):

final answer of \[4x^3\sqrt{2x}\]